# Quantum Error Correction near the Coding Theoretical Bound

### v1

@@ 1 Introduction
First, we present a method for constructing orthogonal protograph matrix pairs. Next, we provide an upper bound on the girth of these protograph matrix pairs. Furthermore, we show that the conditions for achieving this upper bound serve as sufficient and necessary conditions for applying finite field extension method. By replacing the finite field components of the extended protograph matrix pairs with the corresponding companion matrices, we obtain binary orthogonal extended matrix pairs that construct the proposed CSS codes.
@@ 2.2 Decoding method
It is known that any error occurring in a quantum state can be corrected by addressing both bit-flip and phase-flip errors [19]. Binary CSS codes defined by $(H_{X},H_{Z})$ correct errors by estimating the error vector $(\underline{x},\underline{z})$ based on the corresponding syndrome vectors $\underline{s}=H_{X}\underline{x}$ and $\underline{t}=H_{Z}\underline{z}$.
@@ 4 Conclusions
Figure 2: Decoding performance of the proposed code: $f_{m}$ vs. $R$. The plot shows $(f_{m}^{*},R)$, where $f_{m}^{*}$ represents the threshold and $R$ denotes the quantum coding rate. It can be observed that all rates approach the hashing bound.
@@ A.1 Generation of $(\hat{H}_{X},\hat{H}_{Z})$
    This condition ensures333 For $0\leq j,k<J$, $(\hat{H}_{X}\hat{H}^{\mathsf{T}}_{Z})_{j,k}=\sum_{l=0}^{L/2-1}{\bigl{(}F_{-j+l}G_{k-l}+G_{k-l}F_{-j+l}\bigr{)}}=O$. the orthogonality of $\hat{H}_{X}$ and $\hat{H}_{Z}$. Since a cycle is formed along a specific path444 Two paths consist of a sequence of column sub-matrix indices of length $2L$, denoted as $c_{l}^{1}$ and $c_{l}^{2}$ for $l=0,\cdots,L-1$. For row sub-matrices, the traversal alternates vertically. Specifically, $c_{l}^{1}$ is defined as $l$ when $l$ is even and $L-l$ when $l$ is odd, while $c_{l}^{2}$ is defined as $l$ when $l$ is even and $L-l+1$ when $l$ is odd. traversing all sub-matrices, it can be shown that the girth of this construction is upper-bounded by $2L$. The above conditions serve as sufficient conditions for the girth to achieve this upper bound.
@@ C.1 Decoding algorithm
Let $M=PJ$, $N=PL$, and $n=eN$. Using the matrices $H_{X},H_{Z}\in\mathbb{F}_{2}^{eM\times eN}$ and $H_{\Gamma},H_{\Delta}\in\mathbb{F}_{q}^{M\times N}$ provided in the previous sections, we describe the method for performing error correction. The decoding process involves estimating the noise vectors $\underline{x},\underline{z}\in\mathbb{F}_{2}^{eN}$ from the syndromes $\underline{s}=H_{X}\underline{x}$ and $\underline{t}=H_{Z}\underline{z}$, where the matrices $H_{X}$ and $H_{Z}$ are constructed from the non-binary matrices $H_{\Gamma}$ and $H_{\Delta}$ as described in the previous section.
By leveraging the structure of $H_{\Gamma}$ and $H_{\Delta}$ over $\mathbb{F}_{q}$, the proposed method aims to improve decoding performance by interpreting the binary matrices $H_{X}$ and $H_{Z}$ as derived from non-binary protograph-based LDPC codes. For $n=eN$ qubits, the vector representations of X and Z errors are denoted by $\underline{x}$ and $\underline{z}$, respectively. The role of the decoder is to estimate the noise vectors $\underline{x},\underline{z}\in\mathbb{F}_{2}^{eN}$ from the syndromes $\underline{s}=H_{X}\underline{x},\underline{t}=H_{Z}\underline{z}\in\mathbb{F}_{2}^{eM}$.
@@
$p(\underline{x},\underline{z}|\underline{s},\underline{t})$ $\propto\Bigl{(}\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}(H_{X})_{ij}x_{j}=s_{i}\bigr{]}\Bigr{)}\Bigl{(}\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}(H_{Z})_{ij}z_{j}=t_{i}\bigr{]}\Bigr{)}\Bigl{(}\prod_{j=1}^{N}p\left(x_{j},z_{j}\right)\Bigr{)},$ (45)
@@
$\nu_{ij}^{X}(x_{j})$ $=\sum_{(x_{j^{\prime}}):{j^{\prime}\in\partial_{X}(i)}\setminus j}\mathbbm{1}\Bigl{[}\sum_{j^{\prime}\in\partial_{X}(i)}(H_{X})_{ij^{\prime}}x_{j^{\prime}}=s_{i}\Bigr{]}\prod_{j^{\prime}\in\partial_{X}(i)\setminus j}\mu_{j^{\prime}i}^{X}(x_{j^{\prime}}),$ (46)
$\nu_{ij}^{Z}(z_{j})$ $=\sum_{(z_{j^{\prime}}):{j^{\prime}\in\partial_{Z}(i)}\setminus j}\mathbbm{1}\Bigl{[}\sum_{j^{\prime}\in\partial_{Z}(i)}(H_{Z})_{ij^{\prime}}z_{j^{\prime}}=t_{i}\Bigr{]}\prod_{j^{\prime}\in\partial_{Z}(i)\setminus j}\mu_{j^{\prime}i}^{Z}(z_{j^{\prime}}),$ (47)
@@
$\sum_{j\in\partial_{X}(i)}(H_{X})_{ij}\hat{x}_{j}=s_{i},\quad\sum_{j\in\partial_{Z}(i)}(H_{Z})_{ij}\hat{z}_{j}=t_{i},$
@@ C.2 Sum-product algorithm over finite field
$p(\underline{x},\underline{z}|\underline{s},\underline{t})\propto\Bigl{(}\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}\gamma_{ij}\xi_{j}=\sigma_{i}\bigr{]}\Bigr{)}\Bigl{(}\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}\delta_{ij}\zeta_{j}=\tau_{i}\bigr{]}\Bigr{)}\Bigl{(}\prod_{j=1}^{N}p\left(x_{j},z_{j}\right)\Bigr{)}.$ (55)
For $s_{i},t_{i},x_{j},z_{j}\in\mathbb{F}_{2}^{e}$, we define $\sigma_{i},\tau_{i},\xi_{j},\zeta_{j}\in\mathbb{F}_{q}$ are such that $\underline{v}(\sigma_{i})=s_{i},\underline{w}(\tau_{i})=t_{i},\underline{v}(\xi_{j})=x_{j}$ and $\underline{w}(\zeta_{j})=z_{j}$ hold. Note that $\underline{v}(\cdot)$ and $\underline{w}(\cdot)$ were defined in B.
@@
$p(\underline{s}|\underline{x})$ $=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}(H_{X})_{ij}x_{j}=s_{i}\bigr{]}$ (61)
$=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}A(\gamma_{ij})x_{j}=s_{i}\bigr{]}$ (62)
$=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}A(\gamma_{ij})\underline{v}(\xi_{j})=\underline{v}(\sigma_{i})\bigr{]}$ (63)
$=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}\gamma_{ij}\xi_{j}=\sigma_{i}\bigr{]}.$ (64)
For similar reasons, we obtain the following. Note that the companion matrix is given by $A^{\mathsf{T}}$ and the vector representation is given by $\underline{w}(\cdot)$:
$p(\underline{t}|\underline{z})$ $=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}(H_{Z})_{ij}z_{j}=t_{i}\bigr{]}$ (65)
$=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}A^{\mathsf{T}}(\delta_{ij})z_{j}=t_{i}\bigr{]}$ (66)
$=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}A^{\mathsf{T}}(\delta_{ij})\underline{w}(\zeta_{j})=\underline{w}(\tau_{i})\bigr{]}$ (67)

### v2

@@ 1 Introduction
First, we present a method for constructing orthogonal protograph matrix pairs. Next, we provide an upper bound on the girth of sub-graphs of matrix pairs. Furthermore, we show that the conditions for achieving this upper bound serve as sufficient and necessary conditions for applying finite field extension method. By replacing the finite field components of the extended protograph matrix pairs with the corresponding companion matrices, we obtain binary orthogonal extended matrix pairs that construct the proposed CSS codes.
@@ 2.2 Decoding method
It is known that any error occurring in a quantum state can be corrected by addressing both bit-flip and phase-flip errors [19]. Binary CSS codes defined by $(H_{X},H_{Z})$ correct errors by estimating the error vector $(\underline{x},\underline{z})$ based on the corresponding syndrome vectors $\underline{s}=H_{Z}\underline{x}$ and $\underline{t}=H_{X}\underline{z}$.
@@ 4 Conclusions
Figure 2: Decoding performance of the proposed code: $f_{m}^{*}$ vs. $R$. The plot shows $(f_{m}^{*},R)$, where $f_{m}^{*}$ represents the threshold and $R$ denotes the quantum coding rate. It can be observed that all rates approach the hashing bound.
@@ A.1 Construction of $(\hat{H}_{X},\hat{H}_{Z})$
    This condition ensures333 For $0\leq j,k<J$, $(\hat{H}_{X}\hat{H}^{\mathsf{T}}_{Z})_{j,k}=\sum_{l=0}^{L/2-1}{\bigl{(}F_{-j+l}G_{k-l}+G_{k-l}F_{-j+l}\bigr{)}}=O$. the orthogonality of $\hat{H}_{X}$ and $\hat{H}_{Z}$. Since a cycle is formed along a specific path444 Two paths consist of a sequence of column sub-matrix indices of length $2L$, denoted as $c_{l}^{1}$ and $c_{l}^{2}$ for $l=0,\cdots,L-1$. For row sub-matrices, the traversal alternates vertically. Specifically, $c_{l}^{1}$ is defined as $l$ when $l$ is even and $L-l$ when $l$ is odd, while $c_{l}^{2}$ is defined as $l$ when $l$ is even and $L-l+1$ when $l$ is odd. traversing all sub-matrices, it can be shown that the girth of corresponding graph is upper-bounded by $2L$. The above conditions serve as sufficient conditions for the girth to achieve this upper bound.
@@ C.1 Decoding algorithm
Let $M=PJ$, $N=PL$, and $n=eN$. Using the matrices $H_{X},H_{Z}\in\mathbb{F}_{2}^{eM\times eN}$ and $H_{\Gamma},H_{\Delta}\in\mathbb{F}_{q}^{M\times N}$ provided in the previous sections, we describe the method for performing error correction. The decoding process involves estimating the noise vectors $\underline{x},\underline{z}\in\mathbb{F}_{2}^{eN}$ from the syndromes $\underline{s}=H_{Z}\underline{x}$ and $\underline{t}=H_{X}\underline{z}$, where the matrices $H_{X}$ and $H_{Z}$ are constructed from the non-binary matrices $H_{\Gamma}$ and $H_{\Delta}$ as described in the previous section.
By leveraging the structure of $H_{\Gamma}$ and $H_{\Delta}$ over $\mathbb{F}_{q}$, the proposed method aims to improve decoding performance by interpreting the binary matrices $H_{X}$ and $H_{Z}$ as derived from non-binary protograph-based LDPC codes. For $n=eN$ qubits, the vector representations of X and Z errors are denoted by $\underline{x}$ and $\underline{z}$, respectively. The role of the decoder is to estimate the noise vectors $\underline{x},\underline{z}\in\mathbb{F}_{2}^{eN}$ from the syndromes $\underline{s}=H_{Z}\underline{x},\underline{t}=H_{X}\underline{z}\in\mathbb{F}_{2}^{eM}$.
@@
$p(\underline{x},\underline{z}|\underline{s},\underline{t})$ $\propto\Bigl{(}\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(H_{Z})}_{ij}x_{j}=s_{i}\bigr{]}\Bigr{)}\Bigl{(}\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(H_{X})}_{ij}z_{j}=t_{i}\bigr{]}\Bigr{)}\Bigl{(}\prod_{j=1}^{N}p\left(x_{j},z_{j}\right)\Bigr{)},$ (45)
@@
$\nu_{ij}^{X}(x_{j})$ $=\sum_{(x_{j^{\prime}}):{j^{\prime}\in\partial_{X}(i)}\setminus j}\mathbbm{1}\Bigl{[}\sum_{j^{\prime}\in\partial_{X}(i)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(H_{Z})}_{ij^{\prime}}x_{j^{\prime}}=s_{i}\Bigr{]}\prod_{j^{\prime}\in\partial_{X}(i)\setminus j}\mu_{j^{\prime}i}^{X}(x_{j^{\prime}}),$ (46)
$\nu_{ij}^{Z}(z_{j})$ $=\sum_{(z_{j^{\prime}}):{j^{\prime}\in\partial_{Z}(i)}\setminus j}\mathbbm{1}\Bigl{[}\sum_{j^{\prime}\in\partial_{Z}(i)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(H_{X})}_{ij^{\prime}}z_{j^{\prime}}=t_{i}\Bigr{]}\prod_{j^{\prime}\in\partial_{Z}(i)\setminus j}\mu_{j^{\prime}i}^{Z}(z_{j^{\prime}}),$ (47)
@@
$\sum_{j\in\partial_{X}(i)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(H_{Z})}_{ij}\hat{x}_{j}=s_{i},\quad\sum_{j\in\partial_{Z}(i)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(H_{X})}_{ij}\hat{z}_{j}=t_{i},$
@@ C.2 Sum-product algorithm over finite field
$p(\underline{x},\underline{z}|\underline{s},\underline{t})\propto\Bigl{(}\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\delta_{ij}}\xi_{j}=\sigma_{i}\bigr{]}\Bigr{)}\Bigl{(}\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma_{ij}}\zeta_{j}=\tau_{i}\bigr{]}\Bigr{)}\Bigl{(}\prod_{j=1}^{N}p\left(x_{j},z_{j}\right)\Bigr{)}.$ (55)
For $s_{i},t_{i},x_{j},z_{j}\in\mathbb{F}_{2}^{e}$, we define $\sigma_{i},\tau_{i},\xi_{j},\zeta_{j}\in\mathbb{F}_{q}$ are such that $\underline{w}(\sigma_{i})=s_{i},\underline{v}(\tau_{i})=t_{i},\underline{w}(\xi_{j})=x_{j}$ and $\underline{v}(\zeta_{j})=z_{j}$ hold. Note that $\underline{v}(\cdot)$ and $\underline{w}(\cdot)$ were defined in B.
@@
$p(\underline{t}|\underline{z})$ $=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}(H_{X})_{ij}z_{j}=t_{i}\bigr{]}$ (61)
$=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}A(\gamma_{ij})z_{j}=t_{i}\bigr{]}$ (62)
$=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}A(\gamma_{ij})\underline{v}(\zeta_{j})=\underline{v}(\tau_{i})\bigr{]}$ (63)
@@
For similar reasons, we obtain the following. Note that the companion matrix is given by $A$ and the vector representation is given by $\underline{v}(\cdot)$:
$p(\underline{s}|\underline{x})$ $=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}(H_{Z})_{ij}x_{j}=s_{i}\bigr{]}$ (65)
$=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}A^{\mathsf{T}}(\delta_{ij})x_{j}=s_{i}\bigr{]}$ (66)
$=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}A^{\mathsf{T}}(\delta_{ij})\underline{w}(\xi_{j})=\underline{w}(\sigma_{i})\bigr{]}$ (67)
$=\prod_{i=1}^{M}\mathbbm{1}\bigl{[}\sum_{j=1}^{N}\delta_{ij}\xi_{j}=\sigma_{i}\bigr{]}.$ (68)